See this also as dry run for taking beam on Wednesday 10th - full machine closed.
|Time||Phase||Resp||Approx elapsed time|
|Monday 1/Tuesday 2||Dry Run Things should be in a reasonable good shape following previous tests. All circuits required have been available.||OP|
Machine should have been patrolled. Difference this time is CMS should vacate their cavern and patrol.
|Friday pm||Access system - Beam ON|
Final prep of TI8/12, EIS, final beam interlock tests. Time for LBDS arming and sequencer tests with full permit loop (could take longer this time, everything now in the loop).
|Rossano, Bruno/Ben, Etienne||3 hours|
||OP||1 hour checks - trajectory etc.|
||INJ TEAM||1 hour|
|Friday pm||TDI out - threading - beam to IR7 - correct.||Jorg & team||3 hours|
|Late Friday pm||Collimators out - LBDS inject and dump set at >1 turn - thread beam to point 6 - steer into beam dump channel using orbit correctors - first check dump line instrumentation.||OP - BT|
|Late Friday||Dispersion measurements and investigations and measurements for dispersion and matching at the end of TI 8 (trims and rematching if necessary)||Malika & team||8 hours|
|Over night Fri-Sat||Kick-response measurements, optics checks (no preemptive aperture checks)
Take response data across end TI8 and p8 injection region using last corrector in TI8.
|Jorg & team|
Enable Beam Dump Kicker - first try at inject and dump
|Saturday 09:00 - 16:00||
Cryo 8 down - access - recovery
Beam through Alice - screens
SECTOR 23 revisited - check Alice's multiplicity with screens out.
|Evening||Collimator studies point 3||Ralph|
Polarity //TI8 measurements
Foresee access into CMS for vacuum pump-down checks
|Mike, Luca, et al||
|Sunday midday||Cryo point 8 back - possible|
|Sun afternoon||Full machine dry run - pre-cycle, set and trim etc,||OP||6 hours|
|Sunday late - if S78 back||
Commission beam dump region - extract with beam dump kickers - set of measurements with extracted beam - some checks also proposed with beam onto TCDQ (feasible?) Details here
Bren and team
|TI8/ LHC dispersion measurments continued|
|Overnight Sun/Mon||Polarity checks - consolidation if required.||Frank & team||6 hours|
|08:00||Radiaton survey - point 6||Doris Forkel-Wirth|
|08:30||Controls database stop - 20 minute intervention||Ronny Billen|